OpenStudy (anonymous):

Calculate the derivatives of the following real functions by using the chain rule; maybe several times. In each case, determine the inner and outer function to be used, together with their respective areas of definition. a)\(h(x) = sin(cos(x))\)

OpenStudy (zzr0ck3r):

outside first then inside -cos(cos(x))sin(x)

OpenStudy (zzr0ck3r):

understand?

OpenStudy (anonymous):

hmm can you help me more zzrOck? i am noob at mathematics..

OpenStudy (zzr0ck3r):

and your ding calc?

OpenStudy (zzr0ck3r):

well we take the derivative of the outside function d(sin(x))/dx = cos(x) but we leave whats on the inside inside the derivative so cos(cos(x)) then we multiply it by the derivative of the inside function(this is the chain rule) inside function derived is d(cosx)/dx = -sinx so we have -cos(cosx)sinx

OpenStudy (anonymous):

aah ok, and as conclusion we get -cos(cosx)sinx..?

OpenStudy (anonymous):

we dont need to add anything more?

OpenStudy (zzr0ck3r):

correct there is no product rule here if it where (sinx)(cosx) we would

OpenStudy (anonymous):

ok thank you very much zzrOc3r

OpenStudy (zzr0ck3r):

np is it making since? what is the derivative of (3x+2)^2?

OpenStudy (anonymous):

unfortunately not...

OpenStudy (zzr0ck3r):

sorry for spelling I do this on a phone:) and sometimes just assume auto correct got it right:)

OpenStudy (zzr0ck3r):

ok with the example I just gave (3x+2)^2 what is the outside function?

OpenStudy (zzr0ck3r):

have you learned about composite functions?

OpenStudy (anonymous):

np :) hmm i try to understand and calculate

OpenStudy (anonymous):

i learned but it was long ago, and this year i am not going to university, i try to learn at home..

OpenStudy (anonymous):

actually to be honest i am not started to learn, but next month i have exam, very soon i must started to learn it very serious

OpenStudy (zzr0ck3r):

if f(x) = x^2 and g(x) = 3x+2 then, f composite g = f(g(x)) = (3x+2)^2. What we did here was take one function(g) and made it the input of the other function(f) so anyplace I saw a x in x^2 i put (3x+2) this is what is going on with the chaine rule. Our outside function is the funtion that "feeds" on the other. So our outside function on (3x+2)^2 is x^2 and the inside function is 3x+2 so to do d(f(g(x)))/dx = d((3x+2)^2)/dx we take the derivative of the outside function which is x^2 and get 2x, but we leave what is on the inside so we get 2(3x+2) but now we need to use the chaine rule and multiply what we have by the derivative of the inside so we have 2(3x+2)*d(3x+2)/dx = 2(3x+2)*3 = 6(3x+2)

OpenStudy (anonymous):

ok..

OpenStudy (zzr0ck3r):

in your situation let f(x) = sinx and g(x) = cosx so f composite g = f(g(x)) = sin(cosx) then we do the same thing take the dervative of f(x) which is cosx but leave the inside so cos(cosx) and multiply by the derivative of g(x) the inside part so cos(cosx)*d(cosx)/dx = cos(cos(x))*-sinx = -cos(cos(x))sin(x)

OpenStudy (zzr0ck3r):

ok im done if you got any questions just ask:)

OpenStudy (anonymous):

thank you zzrOc3r very much, for detailed explanation, it s really well explained, it helps me much

OpenStudy (anonymous):

wait i look it deeply once more

OpenStudy (zzr0ck3r):

if you want to do another one of your problems just paste it

OpenStudy (anonymous):

zzrOck3r i have some questions but it will take time now, i need to do today 6-7 questions to deliver homework, but in next weeks if you will be here i will catch you to ask detailed questions, its very important for my exam, and exact this kind of question will come in exam

OpenStudy (anonymous):

thank you, but i have other questions to be solved, if you are interested you are welcome

OpenStudy (zzr0ck3r):

np zbrooks@pdx.edu I check it many times a day. I tudor this class and run a lab on it so its pretty fresh:)

OpenStudy (anonymous):

cool, i will send you a message now to your email so that i dont loose your email ;)

OpenStudy (anonymous):

i feel strong for my Exam thank to open study, here so many people who likes to help for math